# Designing Informative Metrics for Few-Shot Example Selection

### v2

@@ Abstract
Pretrained language models (PLMs) have shown remarkable few-shot learning capabilities when provided with properly formatted examples. However, selecting the “best” examples remains an open challenge. We propose a complexity-based prompt selection approach for sequence tagging tasks. This approach avoids the training of a dedicated model for selection of examples, and instead uses certain metrics to align the syntactico-semantic complexity of test sentences and examples. We use both sentence- and word-level metrics to match the complexity of examples to the (test) sentence being considered. Our results demonstrate that our approach extracts greater performance from PLMs: it achieves state-of-the-art performance on few-shot NER, achieving a 5% absolute improvement in F1 score on the CoNLL2003 dataset for GPT-4. We also see large gains of upto 28.85 points (F1/Acc.) in smaller models like GPT-j-6B. [^1] Structured prompting base code taken from https://github.com/blvns/structured_prompting.git.
@@
Rishabh Adiga University of Illinois Urbana-Champaign radiga2@illinois.edu                        Lakshminarayanan Subramanian New York University lakshmi@nyu.edu
Varun Chandrasekaran University of Illinois Urbana-Champaign varunc@illinois.edu
@@ 1 Introduction
Pretrained language models (PLMs) have demonstrated impressive few-shot learning capabilities when prompted with examples for a particular task (Brown et al., 2020). In few-shot prompting, $k$ examples that demonstrate the goal of the downstream task along with the test sample (collectively called the $k$-shot prompt) are provided to the PLM. But the effectiveness of PLMs in few-shot learning depends on the quality of these prompts; this is challenging due to the need for unambiguous instruction (Sclar et al., 2024) which is further complicated due to the task of selecting examples. Appropriately selected examples will ensure that the PLM can generalize effectively and perform optimally across diverse natural language processing (NLP) tasks.
Recent work has shown that specific examples provided in-context can significantly impact the performance of PLMs (Min et al., 2022)Sorensen et al. (2022) notes that specific example combinations lead to better generalization, and connects this with the informativeness of these examples. This suggests that surfacing specific properties of prompts result in improved performance. Most prior work for prompt selection has either trained a prompt retriever to select examples, or used PLMs themselves to label examples (Rubin et al., 2022) to serve as a good prompt. However, both approaches require additional computation and task specific training. Liu et al. (2022) uses nearest neighbours algorithms to select examples. However, this approach solely relies on embedding space properties, and does not consider other metrics related to the label distribution, leading to sub-optimal performance in some tasks.
@@
We conduct our experiments on named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. We show that CP retrieval improves accuracy significantly in all cases.
The contributions of this work are as follows:
  1. 1.
    We propose CP retrieval to select prompts based on sentence and word level complexity metrics (§ 3).
  2. 2.
    We demonstrate improved few-shot accuracy with CP retrieval. Our approach achieves state-of-the-art performance in few-shot NER with a 5% absolute improvement in F1 score on the CoNLL2003 dataset for GPT-4 and upto 28.85 points (F1/Accuracy) in smaller models (§ 5).
@@ 2 Related Work
To perform our experiments, we use the Structured Prompting paradigm introduced by Blevins et al. (2023) as our baseline for named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. Each model that we use is provided $k$ pairs of (sentence, tagged sentence) as the task examples in a $k$-shot prompt. This involves iteratively feeding the PLM a word and its predicted label to get it to label the next word. These are sequence tagging tasks—those that assign labels or tags to tokens or other units of text. Experiments show that PLMs can effectively solve these tasks with just a few demonstrations using this technique.
Liu et al. (2022) has explored strategies for selecting examples to leverage GPT-3’s few-shot learning capabilities. They analyzed the sensitivity of GPT-3’s performance to randomly sampled examples and found significant variability (a phenomenon also explained by Min et al. (2022)).  Liu et al. (2022) proposed a retrieval based approach called KATE that selects semantically similar examples by using $k$-nearest neighbors in the embedding space of pre-trained sentence encoders like BERT and RoBERTa. They found that KATE substantially improved GPT-3’s results over random sampling on natural language understanding and generation tasks. Results with the $k$NN based approach as baseline are demonstrated in Appendix A.2.
@@ 3 Methodology
Our core idea is simple: to select examples that closely resemble the test sample in both semantics and length while ensuring a diverse range of labels for the task. We measure the “complexity” of a candidate example using 3 different sentence-level and word-level metrics: (a) normalized sentence similarity score, (b) normalized smoothed length similarity, and (c) normalized label entropy.
@@
3.1 Normalized Sentence Similarity Score
@@ 3.4 Complexity Score
We use this score to select the $k$ highest scoring train sentences for a $k$-shot prompt. $w_{1}$, $w_{2}$, $w_{3}$ are weights that are set using grid search to optimize scores on the development set for each sequence tagging task. We found that $(w_{1},w_{2},w_{3})$ = $(0.25,0.25,0.5)$ is the best value for NER, $(w_{1},w_{2},w_{3})$ = $(0.2,0.1,0.7)$ is the best value for Chunking and $(w_{1},w_{2},w_{3})$ = $(0.1,0.1,0.8)$ is the best value for POS Tagging. We see that the above set of values indicate that sentence similarity is more important. This is most likely due to the presence of extremely similar examples in the test and training set which are flaws with respect to the dataset.
@@ 4 Experimental Setup
4.1 Tasks
We perform our experiments on 3 different sequence tagging tasks namely NER, POS tagging, and sentence chunking. Each task involves the annotation of tokens in a sequence with specific labels, facilitating the extraction of valuable information and enhancing language understanding. For all datasets, we use a random set of 1000 test samples.
1. NER: This focuses on classifying named entities within a text. The entities can range from persons and organization to locations and dates. For our experiments, we use the widely recognized CoNLL2003 dataset (Tjong Kim Sang and De Meulder, 2003) for measuring the proficiency of our CP retrieval.
@@ 4.1 Tasks
4.2 Models
We perform our experiments on the GPT-Neo series of models which has parameter counts from 125 million to 20 billion (GPT-Neo-125M,GPT-Neo-1.3B, GPT-Neo-2.7B, GPT-j-6B, GPT-NeoX-20B) and additionally 2 black box models (Davinci-002 and GPT-4) for the sequence tagging tasks of NER, POS tagging and sentence chunking.
@@ 4.2 Models
Davinci-002 (Black Box) $20.30$ 40.03 $35.82$ 51.79 46.0446.0446.0446.04333Accuracy has changed with respect to the original paper for this model probably because of the the reasoning in chen2023chatgpts 52.20
@@
Table 1: Results of using complexity score retrieval (CP retrieval) of the best scoring $k$ examples for the few-shot prompt (here $k$ = 5). The baseline is in the structured prompting paradigm (Blevins et al., 2023) with static prompt examples. CP retrieval significantly surpasses the baseline using linguistic structure alone for the prompt (details in Appendix A.1) in all cases. Best values are in boldface. * in the table shows the previously achieved state-of-the-art performance on few-shot NER by Ashok and Lipton (2023).
Figure 1: CP retrieval demonstrates superior performance across all tasks. (a) Performance while varying number of demonstrations $k$ (GPT-Neo-2.7B) for NER. (b) Performance while varying model size for a fixed $k$ = 5 for POS. (c) Performance for the various tasks on GPT-NeoX-20B model with $k$ = 5. Note taht NER and chunking are evaluated based on the F1 score, and POS is evaluated based on accuracy as commonly done.
@@ 5 Results
Our results show that CP retrieval significantly improves few-shot accuracy over the baseline across all models and considered tasks. Table 1 compares the F1 and Accuracy scores (for NER, chunking and POS tagging) for different models while keeping the number of examples constant.
@@
We observe substantial gains using this method with the largest gains on GPT-j-6B which achieves a 28.85 point increase on NER, GPT-Neo-2.7B which achieves a 24.82 point increase in chunking and GPT-Neo-125M which achieves a 15.56 point increase in POS tagging. It is also evident that the gains of example selection drops with increasing model size in Figures 1(b) and 1(c) (with GPT-4 having the least gains) and this is due to the method approaching overall accuracy saturation points. This is also combined with the fact that larger models have seen more diverse training data, so random prompt selections are more than likely to cover the needed distribution and hence, additional gains from optimal selection are diminished. However, even with small gains, we achieve state-of-the-art performance on the CoNLL2003 dataset with a 5% increase over PromptNER methodology (Ashok and Lipton, 2023) for few-shot NER .
@@
From Figure 1(a), we see that our approach enables effective learning (larger gains) with smaller $k$. At $k$ = 2, we see a gain of 19.37 points over the baseline which decreased when $k$ is increased. The percentage gains consistently reduce as we increase the number of examples.
@@
Discussion: While we have focused on complexity metrics for selecting examples in simple prompting setups, what we want to put forth is that this approach can further boost performance when combined with more sophisticated prompting techniques. For example, CP retrieval can be used along with chain-of-thought (CoT) prompting (Wei et al., 2023). It can be used to identify the most information-rich examples for the initial context in CoT prompting which will result in larger gains.
@@ 7 Limitations
Errata
The previous version of this paper contained results pertaining to a dataset involving annotating statements with contextual integrity parameters. This dataset can not be publicly shared, and all related results have been removed (since they will not be reproducible).
Appendix A Appendix
A.1 Example Prompts
@@
A.2 $k$NN Baseline
@@
Table 2: Results of using complexity score retrieval (CP retrieval) of the best scoring $k$ examples for the few-shot prompt (here $k$ = 5). The baseline here is picking the $k$-nearest neighbors + structured prompting paradigm (Blevins et al., 2023).

### v3

@@ Abstract
Pretrained language models (PLMs) have shown remarkable few-shot learning capabilities when provided with properly formatted examples. However, selecting the “best” examples remains an open challenge. We propose a complexity-based prompt selection approach for sequence tagging tasks. This approach avoids the training of a dedicated model for selection of examples, and instead uses certain metrics to align the syntactico-semantic complexity of test sentences and examples. We use both sentence- and word-level metrics to match the complexity of examples to the (test) sentence being considered. Our results demonstrate that our approach extracts greater performance from PLMs: it achieves state-of-the-art performance on few-shot NER, achieving a 5% absolute improvement in F1 score on the CoNLL2003 dataset for GPT-4. We also see large gains of upto 28.85 points (F1/Acc.) in smaller models like GPT-j-6B. [^1] Code and data used for the experiments in this paper is available at https://github.com/RishabhAdiga/Complexity-based-prompt-retrieval.git.
@@
 Rishabh Adiga     Lakshminarayanan Subramanian     Varun Chandrasekaran University of Illinois Urbana-Champaign, IL New York University, NY {radiga2, varunc}@illinois.edu lakshmi@nyu.edu
@@ 1 Introduction
Pretrained language models (PLMs) have demonstrated impressive few-shot learning capabilities when prompted with examples for a particular task (Brown et al., 2020). In few-shot prompting, $k$ examples that demonstrate the goal of the downstream task along with the test sample (collectively called the $k$-shot prompt) are provided to the PLM. However, the effectiveness of PLMs in few-shot learning depends on the “quality” of these prompts; achieving this quality is challenging due to the need for unambiguous instruction (Sclar et al., 2024), which is further complicated by the task of selecting appropriate examples. Appropriately selected examples will ensure that the PLM can generalize effectively and perform well across diverse natural language processing (NLP) tasks.
Recent work has shown that specific examples provided in-context can significantly impact the performance of PLMs (Min et al., 2022)Sorensen et al. (2022) notes that specific example combinations lead to better generalization, and connects this with the informativeness of these examples. This suggests that surfacing “specific” properties of prompts result in improved performance. Most prior work for prompt selection has either trained a prompt retriever to select examples, or used PLMs themselves to label examples (Rubin et al., 2022) to serve as a good prompt. However, both approaches require additional computation and task specific training. Liu et al. (2022) use nearest neighbours algorithms to select examples. However, this approach solely relies on embedding space properties, and does not consider other metrics which model label distribution such as entropy, leading to sub-optimal performance in some tasks.
@@
We conduct our experiments on named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. We also create a new dataset involving tagging of Contextual Integrity (CI) parameters for privacy policies. We show that CP retrieval improves accuracy significantly in all cases.
The contributions of this work are as follows: (1) We propose CP retrieval to select prompts based on sentence and word level complexity metrics (§ 3); and (2) We demonstrate improved few-shot accuracy with CP retrieval. Our approach achieves state-of-the-art performance in few-shot NER with a 5% absolute improvement in F1 score on the CoNLL2003 dataset for GPT-4 and upto 28.85 points (F1/Accuracy) in smaller models (§ 5).
@@ 2 Related Work
To perform our experiments, we use the Structured Prompting paradigm introduced by Blevins et al. (2023) as our baseline for named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. These are sequence tagging tasks—those that assign labels or tags to tokens or other units of text. This involves iteratively feeding the PLM a word and its predicted label to get it to label the next word. PLMs are also provided $k$ pairs of (sentence, tagged sentence) as the task examples in a $k$-shot prompt. Experiments by Brown et al. (2020) show that PLMs can effectively solve these tasks with just a few demonstrations using this technique. We consider another task involving annotating contextual integrity (CI) parameters in privacy policies (Shvartzshnaider et al., 2019) ; we use a generic few-shot prompt as the baseline for this task as shown in Appendix A (Fig. 6).
Liu et al. (2022) have explored strategies for selecting examples to leverage GPT-3’s few-shot learning capabilities. They analyzed the sensitivity of GPT-3’s performance to randomly sampled examples and found significant variability (a phenomenon also explained by Min et al. (2022)).  Liu et al. (2022) proposed a retrieval based approach called KATE that selects semantically similar examples by using $k$-nearest neighbors (in the embedding space of pre-trained sentence encoders like BERT and RoBERTa) which are used as the examples for the $k$-shot prompt. They found that KATE substantially improved GPT-3’s results over random sampling on natural language understanding and generation tasks.
@@ 3 Methodology
Our core idea is simple: to select examples that closely resemble the test sample in both semantics and length while ensuring a diverse range of labels for the task. We measure the “complexity” of a candidate example using 3 different sentence- and word-level metrics: (a) normalized sentence similarity score, (b) normalized smoothed length similarity, and (c) normalized label entropy.
@@
3.1 Normalized Sentence Similarity
@@ 3.4 Complexity Score
We use this score to select the $k$ highest scoring train sentences for a $k$-shot prompt. $w_{1}$, $w_{2}$, $w_{3}$ are weights that are set using grid search to optimize scores on the development set for each sequence tagging task. We found that $(w_{1},w_{2},w_{3})$ = $(0.25,0.25,0.5)$ is the best set for NER, $(w_{1},w_{2},w_{3})$ = $(0.2,0.1,0.7)$ is the best set for Chunking, $(w_{1},w_{2},w_{3})$ = $(0.1,0.1,0.8)$ is the best set for POS Tagging and $(w_{1},w_{2},w_{3})$ = $(0.1,0.1,0.8)$ is the best set for CI tagging. We see that the above set of values indicate that sentence similarity is most important. This is likely due to the presence of extremely similar examples in the test and training set.
@@ 4 Experimental Setup
Tasks: We perform our experiments on 3 different sequence tagging tasks namely NER, POS tagging, and sentence chunking. Each task involves the annotation of tokens in a sequence with specific labels, facilitating the extraction of valuable information and enhancing language understanding. For all datasets, we use a random set of 1000 test samples. For the CI task, we use 100 test samples.
1. NER: This focuses on classifying named entities within a text. The entities can range from persons and organization to locations and dates. For our experiments, we use the widely recognized CoNLL2003 dataset (Tjong Kim Sang and De Meulder, 2003).
@@
4. CI parameters: We additionally created a custom dataset that involves the sequence tagging task of CI parameters for privacy policies. For more details on this dataset, refer to Appendix C.
Models: We perform our experiments on the GPT-Neo series of models which has parameter counts from 125 million to 20 billion (GPT-Neo-125M,GPT-Neo-1.3B, GPT-Neo-2.7B, GPT-j-6B, GPT-NeoX-20B) and additionally 2 black box models (Davinci-002 and GPT-4) for the sequence tagging tasks of NER, POS tagging and sentence chunking. For our custom dataset for CI parameters tagging, we use Llama2 variants.
@@
Davinci-002 (Black Box) $20.30$ 40.03 $35.82$ 51.79 46.0446.0446.0446.04333Accuracy has changed with respect to the original paper for this model probably because of the the reasoning in Chen et al. (2023) 52.20
@@
Table 1: Results of using complexity scores for retrieval of the best $k$ examples (CP retrieval) for the few-shot prompt (here $k$ = 5). The baseline is in the structured prompting paradigm (Blevins et al., 2023) with static prompt examples. CP retrieval significantly surpasses the baseline using linguistic structure alone for the prompt (details in Appendix A) in all cases. Best values are in boldface. * in the table shows the previously achieved state-of-the-art performance on few-shot NER by Ashok and Lipton (2023).
CI tagging (Acc.)
Baseline CP retrieval
Llama2-13B $36.07$ 51.82
Llama2-70B $43.29$ 55.95
Table 2: Results of using complexity score for retrieval (CP retrieval) on the CI parameters tagging task using Llama2 models. Best values are in boldface.
Figure 1: CP retrieval demonstrates superior performance across all tasks. (a) Performance while varying number of demonstrations $k$ (GPT-Neo-2.7B) for NER. (b) Performance while varying model size for a fixed $k$ = 5 for POS. (c) Performance for the various tasks on GPT-NeoX-20B model with $k$ = 5. Note that NER and chunking are evaluated based on the F1 score, and POS is evaluated based on accuracy as commonly done.
@@ 5 Results
Table 1 compares the F1 and Accuracy scores (for NER, chunking and POS tagging) for different models while keeping the number of examples constant. Table 2 provides the same for the CI tagging task on Llama models (performance in a different domain). Our results show that CP retrieval significantly improves few-shot accuracy over the baseline across all models and considered tasks.
@@
We observe substantial gains using our method, with the largest gains on GPT-j-6B which achieves a 28.85 point increase on NER, followed by GPT-Neo-2.7B which achieves a 24.82 point increase in chunking, and finally with GPT-Neo-125M which achieves a 15.56 point increase in POS tagging. It is also evident that the percentage gains of example selection drops with increasing model size in Fig. 1(b) (with GPT-4 having the least gains), and this is due to CP retrieval approaching overall accuracy saturation. Low gains in GPT-4 are also because larger models have seen more diverse training data, so random prompt selections are more than likely to cover the needed distribution. However, even with small gains, we achieve state-of-the-art performance on the CoNLL2003 dataset with a 5% increase over the previous state-of-the-art by Ashok and Lipton (2023) for few-shot NER.
NER (F1) Chunking (F1) POS (Acc.)
Normalized Sentence Similarity $25.39$ $38.16$ $66.82$
Normalized Smoothed Length Similarity $17.22$ $31.50$ $59.91$
Normalized Label Entropy $15.08$ $28.11$ $60.43$
Baseline $12.46$ $24.28$ $55.53$
Table 3: Results of performing ablations to analyze the performance of each individual metric when used for retrieval of $k$ examples for a few-shot prompt (here $k$ = 5).
Figure 2: Normalized confusion matrices for (a) NER (b) Chunking and (c) POS
@@
Impact of $k$: From Fig. 1(a), we see that our approach enables effective learning (larger gains) with smaller $k$. At $k$ = 2, we see a gain of 19.37 points over the baseline which decreased when $k$ is increased. The percentage gains consistently reduce as we increase the number of examples.
@@
Discussion: What we want to put forth is that this approach can further boost performance when combined with more sophisticated prompting techniques. For example, CP retrieval can be used along with chain-of-thought (CoT) prompting (Wei et al., 2023). It can be used to identify the most information-rich examples for the initial context in CoT prompting.
5.1 Ablation Study
Impact of Indvidual Complexity Metrics: Table 3 shows the performance of each individual metric in isolation to retrieve the best $k$ examples for the few-shot prompt. It is evident that all 3 metrics lead to gains in accuracy/F1 scores for each of these datasets. It is also important to observe that the Normalized Sentence Similarity metric leads to the highest gains in performance over the baseline. This is also directly reflected in the optimal values for the weights $(w_{1},w_{2},w_{3})$ in seen in § 3.4 which have a higher value associated with $w_{3}$.
Other Baselines: We also compared our methodology with the $k$NN baseline in Appendix B demonstrating that CP retrieval consistently outperforms $k$NN-based example retrieval. For example, it results in a 4.27% absolute gain in F1 score over the $k$NN baseline for the task of sentence chunking when using GPT-Neo-2.7B model.
Number of Samples: In Appendix D, we show that CP retrieval helps in improving performance even when the number of samples to retrieve from in the train set is reduced. Specifically, when the number of available samples for the NER task is reduced to 1000 (from 11079), CP retrieval achieves an F1 score of 25.98 compared to the baseline random selection score of 12.46. Even with just 30 samples, CP retrieval attains an F1 score of 18.19.
5.2 Error Analysis
In this subsection, we analyze the common errors made by PLMs using our methodology to identify patterns and potential areas for improvement. We do so by looking at the confusion matrices shown in Fig. 2. To plot these, we have zeroed out the diagonal values which represent correct predictions and then perform normalization, making it easier to compare error rates across different classes.
We see that error rate is more evenly spread out for NER (with a max of around 0.25) (c.f. Fig. 2(a)), whereas Chunking (c.f. Fig. 2(b)) and POS (c.f. Fig. 2(c)) have more localized errors. We see vertical bands in the confusion matrices of Chunking and POS, and these vertical bands can indicate systematic biases or recurring mistakes in the model which can be targeted for correction. Specifically, the model over predicts many words as "B-NP" (Beginning of a Noun Phrase) in chunking and incorrectly predicts the "SYM" (Symbol) tag as "PUNCT" (Punctuation) in POS tagging.
@@ 7 Limitations
Appendix
Appendix A Prompts
@@
Context: Classify each word in the the sentence to its contextual integrity parameter (Sender, Attribute, Receiver, None, TP, Subject) Example1: Input Sentence: [We, will, also, disclose, nonpersonally, identifiable, information, to, our, partners, and, other, third, parties, about, how, our, users, collectively, use, the, Sites] Output: [[We, Sender], [will, None], [also, None], [disclose, None], [nonpersonally, Attribute], [identifiable, Attribute], [information, Attribute], [to, None], [our, Receiver], [partners, Receiver], [and, None], [other, Receiver], [third, Receiver], [parties, Receiver], [about, TP], [how, TP], [our, TP], [users, TP], [collectively, TP], [use, TP], [the, TP], [Sites, TP]] Example2: Input Sentence: [If, you, choose, to, use, our, referral, service, to, tell, a, friend, about, Military, or, refer, other, information, on, Military, to, a, friend, we, will, ask, you, for, your, friends, name, and, email, address] Output: [[If, TP], [you, TP], [choose, TP], [to, TP], [use, TP], [our, TP], [referral, TP], [service, TP], [to, TP], [tell, TP], [a, TP], [friend, TP], [about, TP], [Military, Attribute], [or, TP], [refer, TP], [other, Attribute], [information, Attribute], [on, TP], [Military, Attribute], [to, TP], [a, TP], [friend, TP], [we, Receiver], [will, None], [ask, None], [you, Sender], [for, None], [your, Subject], [friends, Subject], [name, Attribute], [and, None], [email, Attribute], [address, Attribute]] Question: Now perform the same task on the input given below and provide output in the same format as above: Input Sentence: [Insert Input]
Figure 6: Prompt structure for the CI tagging task (2-shot prompt shown here)
Appendix B $k$NN Baseline
@@
Table 4: Results of using complexity score retrieval (CP retrieval) of the best scoring $k$ examples for the few-shot prompt (here $k$ = 5). The baseline here is picking the $k$-nearest neighbors + structured prompting paradigm (Blevins et al., 2023). Best values are in boldface which demonstrates that CP retrieval aids in significant performance gains.
Appendix C Contextual Integrity Dataset Details
The Contextual Integrity (CI) parameter tagging task (Shvartzshnaider et al., 2019) involves classifying the words in relevant sentences of privacy policies into the following classes - Sender, Receiver, Subject, Attribute and Transmission Principle. The total counts for each of these classes in the dataset is shown in Table 5. The dataset has a total of 600 samples (sentences) which have been extracted from The OPP-115 Corpus (Online Privacy Policies, set of 115) (Wilson et al., 2016) and labelled for this sequence tagging task. We utilize GPT-4 to create this dataset in a few-shot scenario (the examples for the few-shot prompts are manually labelled). Following this, we obtain the labelled outputs from GPT-4 which is then passed through human verification and relabelling. 3 annotators performed this task such that each sample is verified by 2 annotators and we achieve an average inter-annotator agreement of 0.64 using the popular Cohen’s Kappa metric (McHugh, 2012) on the relabelled classes. We have used 100 samples in the test set and the remaining 500 samples as part of the train set.
Class Frequency
O 4817
SENDER 481
RECEIVER 1058
SUBJECT 509
ATTRIBUTE 3152
TP 6798
Table 5: Total class counts in the CI dataset
Appendix D Dependence on Number of Training Samples Available for Retrieval
While most practical tasks nowadays usually have a large set of examples present to select from, we wanted to check the viability of our methodology as the number of examples for selection reduced. The experiment was performed on GPT-Neo-125M with a 5-shot prompt on the NER datatset. It is evident from Table 6 that reducing the number of examples for selection reduces the F1 score gradually, but there are still significant gains even with a small set of examples to select from when it is compared to the random selection baseline.
No. of samples to select from CP retrieval (F1)
11079 (Entire train data) 29.21
1000 25.98
500 24.38
100 23.74
30 18.19
Baseline-random sample selection (F1)
11079 (Entire train data) 12.46
Table 6: Performance when number of samples to select from (train set) is reduced gradually. Even with just 30 samples to select from, CP retrieval provides a 5.73% absolute gain in F1 score.